Sarah wants her $9000 GIC to double in value in nine years. What annual rate of interest is required for her to reach her goal. Assume interest is compounded annually.
Answer by stanbon(75887)

Sarah wants her $9000 GIC to double in value in nine years. What annual rate of interest is required for her to reach her goal. Assume interest is compounded annually.
-----------------
A(t) = Ao(1+r/n)^(nt)
18000 = 9000(1 + r/1)^(1*9)
2 = (1+r)^9
Take the 9th root of both sides to get:
1+r = 1.08
r = 0.08
------------------
The annual rate must be 8%
==========================
Cheers,
Stan H.